OpenStudy (bridgeni):

What is the value of y so that the ordered pair (5, y) is a solution to the equation -5x + y = -40?

OpenStudy (isaidavila):

(x;y)--->(5;y) X=5 Replace in the equation

OpenStudy (isaidavila):

And find y

OpenStudy (bridgeni):

A. 7 B. -15 C. 9 D. -65

OpenStudy (bridgeni):

Isaid i did that i kept getiing y = 4 idk what im doing wrong

OpenStudy (isaidavila):

Ok -5(5)+y=-40 -25+y=-40 Sum 25 in both sides.... Y=-40+25

OpenStudy (bridgeni):

omg i love you THANK YOU SO Much

OpenStudy (isaidavila):

No problem

OpenStudy (anonymous):

Ok here is a way you find out :) __________________________________________________________ You can do this with any operation really :P okay well if you plug the numbers in, you get.. k2+3=4 then you subtract 3 from 4 and have k2=1 then you divide both sides by 2 and get either k=1/2 or k=0.5 hope that helped. (: